How many milliliters of a 47% alcohol solution must be added to 9 milliliters of a 7% alcholol solution to produce a 17% alcohol solution?
Answer by mananth(16946)

percent ---------------- quantity
Alcohol I 7 ---------------- 9 ml
Alcohol II 47 ---------------- x ml
Mixture 17 ---------------- 9 + x ml

7*9+47x =17(9+x )
63 + 47 x = 153 + 17 x
47 x + -17 x = 153 - -63
30 x = 90
/ 30
x = 3
3 ml of 47 % Alcohol II